20.0 g of sodium metal is dropped into a beaker of water to produce hydrogen gas and

sodium hydroxide. What volume of hydrogen gas will be produced if the temperature is 30oC and the pressure is 125 kPa? (Be sure to provide a balanced chemical equation with

your solution.)


1
Expert's answer
2021-03-31T08:05:53-0400

2H2O + 2Na = 2NaOH + H2

Moles of Na = 20/23

0.87moles

Na:H2

2:1

Moles of H2= 0.87÷2

0.435moles

Pv=nRT

125× v= 0.435 × 8.314 × 303

Volume of hydrogen = 8.766